OpenStudy (anonymous):

Given f(x)=5x+square root 7, compute f(x)-f(-2)/x-(-2). Assume x does not equal to -2. Show all work. I do not understand this function equation, as I am used to having 0's as something not equal to. Help is greatly appreciated.

OpenStudy (richyw):

\[f(x)=5x+\sqrt{7}\]\[f(-2)=-10+\sqrt{7}\]\[\frac{f(x)-f(-2)}{x-(-2)}=\frac{f(x)+10-\sqrt{7}}{x+2}\]

OpenStudy (kc_kennylau):

\[\Large\frac{f(x)-f(-2)}{x-(-2)}=\frac{5x+\sqrt7+10-\sqrt7}{x+2}=\frac{5x+10}{x+2}=5\]

OpenStudy (anonymous):

thanks guys for your help!